OpenStudy (anonymous):

evaluate the definite integral [0,a] xsqrt(x^2+a^2)dx (a>0)

OpenStudy (anonymous):

x^2 + a^2 = t use this :)

OpenStudy (anonymous):

still not quite sure what to do

OpenStudy (anonymous):

Hmm substitution

OpenStudy (anonymous):

x^2 + a^2 = t 2x = dt/dx dx = dt/2x

OpenStudy (anonymous):

now you can do it just put it on paper

OpenStudy (anonymous):

you awarded me my 3,500th medal :)